OpenStudy (anonymous):

When a coin is flipped TWICE, what is the chance of getting a head AND a tail?

OpenStudy (anonymous):

In what terms do you need the answer? Ratio? Percentage? Each time you flip the coin you have a 50/50 chance of a head and a 50/50 chance of a tail....

OpenStudy (anonymous):

Probably in ratio terms rather than percentage.

thomaster (thomaster):

the first time you flip the coin, you have a 50% chance, or 1/2. The next flip you have a 50% chance again, so another 1/2 \(\sf\dfrac{1}2*\dfrac{1}2=\dfrac{1}4\) (or 25%)

OpenStudy (anonymous):

Thank you for your help!